cost sheet for Largo Corp, showed the following material cost related to one unit of produet: 4pieces at $4 per piece $16 per unit During February 19X1,10,000 units were produced.The costs recorded showed the following Purchase of material 8000 pieces at $3.80 Issue of material 43,000 pieces Determine the material Give the significance and possible explanation of each materials variance What are some causes of unfavorable materials purchase price variances? The purchasing department of a large corporation using the responsibility accounting system purchase price variance and material quantity (usage) variance. delays placing an assembly department order for 10,000 parts at $2.00 each. By the time the additional charges incurred? Why? The standard cost sheet lor York Corp. shows the following cost for a unit of product price has risen to S220 Which department should be assigned the Direct labor: 2% hours at $7 an hour During February 19X1, 10,000 units were completed. The actual cost incurred was 28,000 hours at $8 per hour-$224,000 Determine the labor rate variance and labor efficiency variance . Give the significance and possible explanation of each labor variance Direct labor and variable overhead standards per finished unit for Century-Fox Metals Company are as follow: variable overhead, 10 hours at $2.00 per hour. During October,5,000 units were produced. Direct labor hours used were 52,000 bours Actual variable overbead costs were $109,200. Deterinine the spending variance and efficiency variance for variable overhead. Give the significance and possible explanation of each variable overbead variance E. T. Toys Inc. produces toys for national distibution. The management has recently established a standard cost system to control costs The standards on a particular toy are Materials Labor: 12 pieces per toy at $0.56 per piece 2 hours per toy at $3.75 per hour During the month of Fiebruary 19A, the company produced 1,000 toys Production data for the month are as follows Materials: 17500 pieces were purchased for use in production, at a toral cost of $8,925 of which 3,500 pieces were still in inventory at the end of the month - Labor: 2,500 hours were worked, at a cost of $10.500. (a) Calculate the materials purchase price variance b) Calculate the materials quantity (usage) variance (c) Calculate the labor rate variance (d) Cakulate the labor efficiency variance
Answer #1

Material Purchase Price variance = (Standard Price – Actual Price)*Actual Quantity Purchased

= (4-3.80)*48,000

= 9,600 F

Material Usage Variance = (Standard Quantity – Actual Quantity)*Standard Price

= (10,000*4 – 43,000)*4

= 12,000 U

Material Purchase Price Variance is favorable which indicates that the actual price paid for material was less than the standard price set which is good for the company.

Possible Reasons are :

1.Change in Market Rate

2.Purchase of Lower Quality Material

Unfavorable Material Usage Variance indicates that the material used in the process is more than the standard quantity of material which should have been used.

Possible Reasons are:

1.Lower quality of material

2.Inefficient Labour

Labor rate Variance = (Standard Rate – Actual Rate)*Actual Hours

= (7-8)*28,000

= 28,000 U

Labor Efficiency Variance = (Standard Hours – Actual Hours)*Standard rate

= (2.75*10,000 – 28,000)*7

= 3,500 U

Both variances are unfavourable which indicates that the labor is paid at more than the standard rate and labor is inefficient as well i.e. it makes more than the standard time set for production.

a)Material Price Variance = (Standard Price – Actual Price)*Actual Quantity

= (0.56 – Actual Price)*17,500

= 9,800 – 8,925

= 875 F

Material Usage Variance = (Standard Quantity – Actual Quantity)*Standard Price

= (12*1,000 – 14,000)*0.56

= 1,120 U

Labor rate Variance = (Standard rate – Actual rate)*Actual Hours

= (3.75 – actual rate)*2,500

= 9,375 – 10,500

= 1,125 U

Labor Efficiency Variance = (Standard Hours – Actual Hours)*Standard Rate

= (2*1,000 – 2,500)*3.75

= 1,875 U